2(n+4)=8-6(n-4)
2n+8=8 -6n+24
-6n+2n=-4n 8+24
-4n + 32
-4n/4=32/4
n=8
NOT ON MY ANSWER SHEET
A. 0
B. 16
C. 3
D. -3

Answer by ewatrrr(24785)

 
Hi,
2(n+4)=8-6(n-4)
2n+8 = 8 -6n+24 |Good work eliminating the parenthesis
8n = 24 |Adding 6n and subtracting 8 from both sides of the Equation
n = 3